# FPGA-Enabled Machine Learning Applications in Earth Observation: A Systematic Review

### v1

@@ 6.3. Recommendations for Future Work
Remote Sensing Good Practices
We encourage researchers to use open-source data or to publish their datasets. When a full release is not possible due to confidentiality agreements, studies should detail, as rigorously as possible, sensor specifications, selected scenes, ground truth, data splits, etc. Such transparency is crucial for future readers to fully comprehend the problem and replicate results.
Comprehensive Reporting of Metrics
A significant limitation across many reviewed studies is the lack of thorough metric reporting, which prevents comprehensive evaluation and comparison. Fundamental performance indicators are often incomplete. For example, object detection works may report mAP, but omit mIoU. Similarly, hardware results may mention throughput and power, but use inconsistent bit-widths (e.g., i4 vs. fp32), skewing energy efficiency comparisons. Even more critical, memory usage is rarely analyzed in depth. While some papers mention memory footprint, very few examine off-chip memory transfers—a common bottleneck for FPGA designs (Abdelouahab et al., 2018; Cong et al., 2018). We also want to highlight the recurrent confusion around the terms GOP, GOPs, and GOP/s. We recommend using GOP to denote computational complexity and GOP/s for computational throughput. Similarly, the term ”throughput” is often used without clarification, though it may refer to either computational throughput [GOP/s] or frame throughput [FPS].
Micro-architecture of Convolution Kernels
Despite the ubiquity of convolution operations in the surveyed applications, most HDL and HLS studies re-implement convolution kernels from scratch, often without an in-depth study or comparison to established designs. This is surprising given the rich body of prior work on optimized CNN accelerators for FPGAs (Abdelouahab et al., 2018; Wang et al., 2019; Abdelfattah et al., 2018; Umuroglu et al., 2017), as well as mature ASIC tape-outs from both academia (Chen et al., 2019) and industry (Jouppi et al., 2021). Rather than reinventing the convolution micro-architecture, new studies could build on these foundations and shift focus toward application-specific challenges. Furthermore, we have observed that many studies highlight low FPGA resource utilization as a positive outcome, however, it may also indicate poor scalability. Ideally, FPGA resources should be fully utilized, except when limited by external factors like power budgets or memory bandwidth.
Code Availability
Among the analyzed articles, only few provide accessible code repositories. While we recognize the difficulty of sharing runnable HDL environments, recent tools like Vitis AI (AMD, 2019)—which runs in containers and through customized Python scripts—facilitate code sharing. We strongly encourage future authors to share their code and scripts, as they provide valuable information for the readers and increase the long-term impact of their work.

### v2

@@ 6.3. Recommendations for Future Work
We encourage researchers to use open-source data or to publish their datasets. When a full release is not possible due to confidentiality agreements, studies should detail, as rigorously as possible, sensor specifications, selected scenes, ground truth, data splits, etc. Such transparency is crucial for future readers to fully comprehend the problem and replicate results. In addition, we strongly encourage future authors to share their code and scripts as thoroughly as possible, as this provides valuable information for the readers and increases the long-term impact of their work. Furthermore, a significant limitation across many reviewed studies is the lack of thorough metric reporting, which prevents comprehensive evaluation and comparison. Fundamental performance indicators are often incomplete. For example, some object detection works may report mAP but omit mIoU, preventing comparisons across studies. Similarly, hardware results may mention throughput and power, but use inconsistent bit-widths (e.g., i4 vs. fp32), skewing energy efficiency comparisons. Even more critical, memory usage is rarely analyzed in depth. While some papers mention model footprint, very few examine off-chip memory transfers—a common bottleneck for FPGA designs (Abdelouahab et al., 2018; Cong et al., 2018).
Future research should also build upon existing work in FPGA implementations. Despite the ubiquity of convolution operations in the surveyed applications, most HDL and HLS studies re-implement convolution kernels from scratch, often without an in-depth study or comparison to established designs. This is surprising given the rich body of prior work on optimized CNN accelerators for FPGAs (Abdelouahab et al., 2018; Wang et al., 2019; Abdelfattah et al., 2018; Umuroglu et al., 2017), as well as mature ASIC tape-outs from both academia (Chen et al., 2019) and industry (Jouppi et al., 2023). Rather than reinventing the convolutional microarchitecture, new studies could build on these foundations and shift their focus toward application-specific challenges. Furthermore, we have observed that many studies highlight low FPGA resource utilization as a positive outcome; however, it may also indicate poor scalability. Ideally, FPGA resources should be fully utilized, except when limited by external factors like power budgets or memory bandwidth.